OpenStudy (anonymous):

Mr gomez rode to work averaging 45 mi per hour. he averaged 55 MPH on the trip home from work. if he spent a total of 6 hours driving to and from work, how many miles is his work from his home?

OpenStudy (asnaseer):

lets say his work is x miles away. then, if he travelled an average of 45 mile per hour going to work, it would take him x/45 hours. similarly, if he travelled an average of 55 miles per hour coming back home, it would take him x/55 hours. so we can write:\[\frac{x}{45}+\frac{x}{55}=6\]solve this to find x.

OpenStudy (anonymous):

do we cross Multiply?

OpenStudy (asnaseer):

do you know how to add fractions?

OpenStudy (anonymous):

i dont rember no

OpenStudy (asnaseer):

ok, first you need to find the lowest common multiple (or LCM) of 45 and 55. do you know what that is?

OpenStudy (asnaseer):

lets start with a simpler fraction problem. lets say we had:\[\frac{1}{3}+\frac{1}{4}\]the first thing we need is the LCM of 3 and 4, do you know what that is?

OpenStudy (anonymous):

1?...

OpenStudy (asnaseer):

no - you need to find the lowest number that is both "multiple" of 3 and a multiple of "4"

OpenStudy (anonymous):

2?

OpenStudy (asnaseer):

e.g.: 3 x 1 = 3 <-- reject as this is NOT a multiple of 4 3 x 2 = 6 <-- reject as this is NOT a multiple of 4 3 x 3 = 9 <-- reject as this is NOT a multiple of 4 3 x 4 = 12 <-- BINGO! this is ALSO a multiple of 4 so 12 is the LCM of 3 and 4 understand?

OpenStudy (anonymous):

oh okay... lol

OpenStudy (anonymous):

so 495?

OpenStudy (asnaseer):

so next we can re-write the simpler fraction as:\[\frac{1}{3}+\frac{1}{4}=\frac{1}{3}*\frac{4}{4}+\frac{1}{4}*\frac{3}{3}=\frac{4}{12}+\frac{3}{12}=\frac{7}{12}\]

OpenStudy (asnaseer):

yes - 495 is the LCM of 45 and 55 - good

OpenStudy (asnaseer):

so do you think you can do the question now?

OpenStudy (anonymous):

yep ive got it now ty

OpenStudy (asnaseer):

yw